Can anyone please help me with a few questions?
81^3x-3=729^*+2 find the solution set

Answer by fractalier(6550)

From
81^(3x-3)=729^(x+2)
we change to powers of nine and get
9^2^(3x-3) = 9^3^(x+2)
9^(6x - 6) = 9^(3x + 6)
If the bases are equal, the exponents are too, so we get
6x - 6 = 3x + 6
3x = 12
x = 4